OpenStudy (inkyvoyd):

Integrate \(\frac{-\frac{1}{\sqrt{2}v}}{v^2+\sqrt{2}v+1}+\frac{\frac{1}{\sqrt{2}}v}{v^2-\sqrt{2}v+1}\)

OpenStudy (inkyvoyd):

@nbouscal will know what I am talking about, or maybe @Limitless

OpenStudy (inkyvoyd):

Wait, correction.

OpenStudy (inkyvoyd):

\(\frac{-\frac{1}{\sqrt{2}}v}{v^2+\sqrt{2}v+1}+\frac{\frac{1}{\sqrt{2}}v}{v^2-\sqrt{2}v+1}\)

OpenStudy (inkyvoyd):

Am I supposed to complete the square?

OpenStudy (inkyvoyd):

@dpaInc , pwease help me!

OpenStudy (anonymous):

can't see it so i'm gonna copy it... and make it huge... \[\huge \frac{-\frac{1}{\sqrt{2}}v}{v^2+\sqrt{2}v+1}+\frac{\frac{1}{\sqrt{2}}v}{v^2-\sqrt{2}v+1} \]

OpenStudy (inkyvoyd):

@dpaInc , how do you copy it like that?

OpenStudy (anonymous):

right click-->show math as-->tex commands

OpenStudy (anonymous):

\[\huge \int\frac{-\frac{1}{\sqrt{2}}v}{v^2+\sqrt{2}v+1}+\frac{\frac{1}{\sqrt{2}}v}{v^2-\sqrt{2}v+1}dv \]

OpenStudy (inkyvoyd):

Alright, now solve NAO!

OpenStudy (inkyvoyd):

xD

OpenStudy (inkyvoyd):

btw, this is the equation I got after trying 3 substitutions and seperating the fraction.

OpenStudy (anonymous):

\[\large \int\frac{-\frac{1}{\sqrt{2}}v}{v^2+\sqrt{2}v+1}+\frac{\frac{1}{\sqrt{2}}v}{v^2-\sqrt{2}v+1}dv\] \[\large \frac{-1}{\sqrt2}\int\frac{v}{v^2+\sqrt{2}v+1}dv+\frac{1}{\sqrt2}\int \frac{v}{v^2-\sqrt{2}v+1}dv\] maybe i should just draw...

OpenStudy (inkyvoyd):

lolol

OpenStudy (inkyvoyd):

@dpaInc , now do I complete the square?

OpenStudy (anonymous):

yeah... that's what i was about to do... but i think there is a formula... hang on....

OpenStudy (anonymous):

there it is.... #13....:) http://integral-table.com/downloads/single-page-integral-table.pdf

OpenStudy (anonymous):

i mean #16...l

OpenStudy (inkyvoyd):

\(\huge \frac{-1}{\sqrt2}\int\frac{v}{(v+\frac{\sqrt{2}}{2})^2+\frac{1}{2}}dv+\frac{1}{\sqrt2}\int \frac{v}{(v-\frac{\sqrt{2}}{2})^2+\frac{1}{2}}dv\)

OpenStudy (inkyvoyd):

But isn't using tables cheating?

OpenStudy (anonymous):

no... of course not...

OpenStudy (inkyvoyd):

But... I don't know where the tables come from... >.<

OpenStudy (anonymous):

it's like using all the different tires for your car... you're not going to make/invent your own tires are you? oh wait.. do you drive yet?

OpenStudy (inkyvoyd):

No lol. I'M GOING TO BUILD MY CAR FROM SCRATCH MUAHAHAHA

OpenStudy (anonymous):

i don't think it's cheating... it's a good shortcut to cut through all that stuff..

OpenStudy (inkyvoyd):

Okay... @dpaInc , without showing me the math, can you tell me what to do after I have completed the square?

OpenStudy (anonymous):

and as for me, i haven't really started using the integral tables until i saw everyon virtually used them here on OS.... i actually went through proving stuff because I don't use/memorize the tables...

OpenStudy (inkyvoyd):

@dpaInc , my problem is that I'm not sure how to get the results shown in the tables.

OpenStudy (anonymous):

let me complete the square on the first integral... i haven't done it for a while and i guess i need the practice...

OpenStudy (inkyvoyd):

I already did in my last tex post :)

OpenStudy (inkyvoyd):

\huge \frac{-1}{\sqrt2}\int\frac{v}{(v+\frac{\sqrt{2}}{2})^2+\frac{1}{2}}dv+\frac{1}{\sqrt2}\int \frac{v}{(v-\frac{\sqrt{2}}{2})^2+\frac{1}{2}}dv

OpenStudy (anonymous):

oh... i see you did it already....

OpenStudy (inkyvoyd):

xD

OpenStudy (anonymous):

you are solving the problem in a wrong way

OpenStudy (inkyvoyd):

@nitz , what am I doing wrong?

OpenStudy (anonymous):

see firstly make numerator as the derivative of denominator

OpenStudy (inkyvoyd):

Can you show me?

OpenStudy (anonymous):

i am talking about 1st integration ie before addition

OpenStudy (inkyvoyd):

?

OpenStudy (anonymous):

in it, firstly multiply the numerator with 2 and add and subtract \[\sqrt{ 2} \] in it

OpenStudy (inkyvoyd):

You mean at the very start of the problem?

OpenStudy (anonymous):

ya

OpenStudy (inkyvoyd):

but but but

OpenStudy (inkyvoyd):

\(\huge \frac{-1}{\sqrt2}\int\frac{v}{(v+\frac{\sqrt{2}}{2})^2+\frac{1}{2}}dv+\frac{1}{\sqrt2}\int \frac{v}{(v-\frac{\sqrt{2}}{2})^2+\frac{1}{2}}dv\) we already got to here >.<

OpenStudy (anonymous):

you cant solve it further in this case

OpenStudy (anonymous):

these are direct problems

OpenStudy (anonymous):

these are direct solutions

OpenStudy (anonymous):

you can apply direct formula

OpenStudy (anonymous):

i think it works....